nini:

@shadow A department store buys 400 shirts at a cost of ​$9600 and sells them at a selling price of ​$ each. Find the percent markup

Shadow:

What is the selling price

nini:

oh srry $ 30

Shadow:

First do 9600/400 to find the cost per shirt

nini:

24

Shadow:

Then we can say 24x = 30 Since they're selling it for $30, and their cost per shirt was 24.

Shadow:

Solving for x is simple as dividing 24 from both sides x = 30/24 x = 1.25

Shadow:

So there was an increase of 25%, since 125% of 24 is 30.

nini:

?

Shadow:

This math works out since 25% of 24 is 6, and 24 + 6 = 30

nini:

so its 30

Shadow:

So the markup % was 25%